OpenStudy (anonymous):

The circle below is centered at the point (8, 4) and has a radius of length 4. What is its equation?

OpenStudy (anonymous):

The standard form is \[(x-h)^2+ (y-k)^2 = r^2\] where (h,k) are the coordinates of the center of the circle and r is the origin. You should be able to do it from there

OpenStudy (anonymous):

Sorry r is the radius

OpenStudy (anonymous):

is it (x + 8)2 + (y - 4)2 = 16?

OpenStudy (anonymous):

or.....(x - 8)2 + (y - 4)2 = 16

OpenStudy (anonymous):

Yeah, the second one. Nice!

OpenStudy (anonymous):

thanks again(:

OpenStudy (anonymous):

No problem :)